OpenStudy (anonymous):

-4+3i someone help please

OpenStudy (anonymous):

...what's the question?

OpenStudy (anonymous):

\[-4+\sqrt{-3}\]

OpenStudy (anonymous):

That's wrong. It's -4 + the square root of negative 9.

OpenStudy (anonymous):

for magnitude on the argand diagram is = \[\sqrt{(-4)^{2}+(3\sqrt{-1})^{2}}\]=\[\sqrt{7}\]